THE CHARACTERIZATION OF HYPERFINE EFFECTS IN YTTERBIUM HALIDE DIATOMIC MOLECULES USING MWFT SPECTROSCOPY

\underline{CAMERON S. DICKINSON}, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia, B3H 4J3, CANADA; NICHOLAS R. WALKER, Department of Chemistry, University of British Columbia, Vancouver, British Columbia, V6T 1Z1, CANADA; JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia, B3H 4J3, CANADA; MICHAEL C. L. GERRY, Department of Chemistry, University of British Columbia, Vancouver, British Columbia, V6T 1Z1, CANADA.

The Ytterbium atom is isovalent with the alkaline earth metals by virtue of its 4f^{14}6s^{2} ground state configuration. Data obtained from studies on diatomic molecules containing Yb therefore provide appropriate parameters for comparison with alkaline earth metal analogs. In this work, the microwave spectra of YbBr, YbCl and YbF have been assigned for the 2Σ^{+}, v = 0 ground state. Precise rotational constants and several hyperfine parameters have been obtained. In addition, studies of the v = 1 levels for the most abundant isotopomers of YbF and YbCl have provided equilibrium bond lengths and a measure of the vibrational dependence for each of the fitted parameters. The determined hyperfine constants have been used to provide information about the electronic structures of the molecules.